If a Willy bat flies from x,y,z coordinate (-2m,9m,-3m) to coordinate (6m,-2m,-3m).what is the displacement change in r to n unit vector notation is ∆r parallel to one of the three coordinate plane. If so, which plane(vectors)
1
Expert's answer
2019-08-22T09:28:32-0400

Displacement is the change of position of a particle with respect to an origin, that is:


"\Delta \vec{r}=\vec{r}_{f}-\vec{r}_{i}"


Where:

  • Initial position "\vec{r}_{i}=(-2,9,-3)m"
  • Final position "\vec{r}_{f}=(6,-2,-3)m"


Evaluating numerically.


"\Delta \vec{r}=(6,-2,-3)m-(-2,9,-3)m"


"\Delta \vec{r}=(6+2,-2-9,-3+3)m"


"\Delta \vec{r}=(8,-11,0)m"


The zero component (Z) indicates that there was no displacement on that axis, therefore, the displacement was parallel to the plane formed by the other two components, in this case.


The displacement is parallel to the X-Y plane